A diver makes 2.5 complete revolutions on the way from a 9.4 m high platform to the water. Assuming zero initial vertical velocity, find the diver's average angular velocity during a dive.
rad/s
Answer #1

Time t is taken to reach water

Using equation

s = ut+0.5gt2

u = 0 , s = 9.4 m, g = 9.8 m/s2

=> t = sqrt(2*9.4/9.8) = 1.385 sec

average angular velocity = 2.5*2*pi/1.385 = 11.346 rad/sec